OpenStudy (anonymous):

Solve : x^2-y=7 x^2+y^2= 49

OpenStudy (nory):

Try multiplying one equation by -1, then adding the two equations.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so it be -x^2+y=-7?

OpenStudy (nory):

Yes. Now add that to the second equation. What do you get?

OpenStudy (anonymous):

y+y^2=42?

OpenStudy (nory):

Yes! Now this is just a quadratic equation. Get all terms on one side and solve it. What do you get?

OpenStudy (anonymous):

y^2+y=-42

OpenStudy (nory):

How did that 42 get negative? The answer is y^2 + y - 42 = 0. Now solve it. What do you get?

OpenStudy (anonymous):

positive 42 lol my bad

OpenStudy (anonymous):

(y+7) (y-6)?

OpenStudy (nory):

Good. Now you know how to solve that.

OpenStudy (anonymous):

thats it?

OpenStudy (nory):

Indeed it is. Don't forget to plug your value of y in to one of the original equations to get x when you're done! Good work. A medal's coming your way. :)

OpenStudy (anonymous):

thanks! but which solution do i use? both? so i have 2 different ordered pairs?